Please explain this for me. The perimeter of a rectangular billboard is 60 feet and its length is 6 feet longer than its width. Find the dimensions of the billboard.
Answer by mananth(12270)

Perimeter = 60 feet
Width = x ft
Length ( x + 8 )

Perimeter =2(L+W)
Perimeter = 2((x + 8)+ x)
Perimeter = 2((x+8+ x)
Perimeter = 16 + 2 x
60 = 2 x + 16
60 - = 2 x
60 -16 = 2 x
44 = 2 x
x= 22 ft Width
Length = 30